# Unveiling Topological Structures in Text: A Comprehensive Survey of Topological Data Analysis Applications in NLP

### v2

@@ 5.1 LLMs to Convert TDA Concepts to Codes
One of the significant challenges in applying TDA to NLP tasks is the steep learning curve associated with its mathematical foundations, which are often accessible only to expert audiences. Moreover, the theorists who develop and understand these advanced concepts do not always collaborate with computational scientists to translate them into executable code. To address this gap, Liu et al. (2024) propose leveraging ChatGPT to generate Python code for TDA concepts by training it on these mathematical foundations. Their findings suggest that ChatGPT can alleviate this bottleneck, particularly for complex TDA concepts like hypergraphs, digraphs, and persistent harmonic space, which have not been as heavily explored as the vietoris-rips complex Liu et al. (2024). Similarly, experts can develop specialized code generators, such as fine-tuning models like Code-Llama Roziere et al. (2023) on TDA concepts. Creating a dedicated LLM for TDA code generation could significantly lower the entry barrier, encouraging the NLP community to explore TDA applications more innovatively.
@@ 5.2 Theoretical Approaches Connecting TDA Features with Linguistic Principles
There is a need for theoretical approaches that better tie TDA features to linguistic phenomena. For instance, Draganov and Skiena (2024) investigate the shape of words and their embeddings in Indo-European languages and finds similar conclusions as Port et al. (2018) and Port et al. (2022), who investigate the syntactic topological space of such languages. They found that TDA features represent historical facts as in languages clustered closely together are similar or influenced by the other. Finally, we observe only 11 theoretical TDA works in NLP, compared to the 70+ non-theoretical ones. Thus, we need more theoretical TDA approaches as it is impractical to achieve the depth and understanding of performance from a topological perspective without more investigations.
@@ 6 Conclusion
Our world is currently experiencing an explosion of data and an equally explosion of computational techniques to process such data. Machine Learning (ML) is the most popular of these computational methods, however, while its benefits are numerous, it has a few limitations. The biggest of the limitations of ML is its inability to sufficiently process data that is high-dimensional, imbalanced, noisy, and scarce. Therefore, a small community of NLP researchers emerged to tackle this limitation by proposing using TDA to tackle difficult NLP tasks. These researchers employ two TDA techniques - Persistent Homology and Mapper to solve NLP tasks using theoretical and non-theoretical approaches. This yielded 87 papers which we comprehensively surveyed in this paper. Finally, we conclude that while the applications of TDA in NLP have come a long way from 2013 to now, there is still room for improvement, specifically in reducing the barrier of entry for non-TDA experts to apply it to their NLP tasks.
Appendix A Appendix
@@
Fitz et al. (2024) Topological “Shape” of Words Measure of topological complexity of Transformer-based hidden representations Persistent Homology
@@
Proskura and Zaytsev (2024a) Ensemble model selection Pre-trained (Transformers - CLS) Unsupervised Text Persistent Homology
@@
Proskura and Zaytsev (2024b) estimation of weights for ensembles of classification models Pre-trained (Transformers - Attention, CLS) Supervised Text Persistent Homology

### v3

@@ 6 Open Problems and Future Directions
6.1 Pros and Cons of TDA in NLP
There are several benefits and challenges to adopting TDA techniques to solve non-trivial problems. We outline these pros and cons for NLP problems in Table 3. The most glaring benefit and challenge are that TDA is effective in low-resource settings and the high computational cost of employing TDA, respectively. This benefit makes TDA particularly attractive for problems with noisy, insufficient, heterogeneous, and high-dimensional datasets. Although the high computational cost presents challenges for the adoption of TDA, recent improvements in computational resources mitigate this challenge. Finally, below we discuss how researchers can leverage these benefits to address the open problems and future directions.
@@ 6.2 LLMs to Convert TDA Concepts to Codes
One of the major challenges in applying TDA to NLP tasks is the steep learning curve associated with its mathematical foundations, which are often accessible only to expert audiences. Moreover, the theorists who develop and understand these advanced concepts do not always collaborate with computational scientists to translate them into executable code. To address this gap, Liu et al. (2024) proposes leveraging ChatGPT to generate Python code for TDA concepts by training it on these mathematical foundations. Their findings suggest that ChatGPT can alleviate this bottleneck, particularly for complex TDA concepts like hypergraphs, digraphs, and persistent harmonic space, which have not been as heavily explored as the Vietoris-Rips complex Liu et al. (2024). Similarly, experts can develop specialized code generators, such as fine-tuning models like Code-Llama Roziere et al. (2023) on TDA concepts. Creating a dedicated LLM for TDA code generation could significantly lower the entry barrier, encouraging the NLP community to explore TDA applications more innovatively.
@@ 6.3 Theoretical Approaches Connecting TDA Features with Linguistic Principles
There is a need for theoretical approaches that better tie TDA features to linguistic phenomena. For instance, Draganov and Skiena (2024) investigate the shape of words and their embeddings in Indo-European languages and find similar conclusions as Port et al. (2018) and Port et al. (2022), who investigate the syntactic topological space of such languages. They found that TDA features represent historical facts as in languages clustered closely together are similar or influenced by the other. Finally, we observe only 12 theoretical TDA works in NLP, compared to over 80 non-theoretical ones. Thus, we need more theoretical TDA approaches as the depth and understanding of performance from a topological perspective without more investigations.
@@ 7 Conclusion
Our world is currently experiencing an explosion of data and an equally explosion of computational techniques to process such data. Machine Learning (ML) is the most popular of these computational methods; however, while its benefits are numerous, it has a few limitations. The biggest of the limitations of ML is its inability to sufficiently process data that is high-dimensional, imbalanced, noisy, and scarce. Therefore, a small community of NLP researchers emerged to tackle this limitation by proposing using TDA to tackle difficult NLP tasks. These researchers employ two TDA techniques - Persistent Homology and Mapper to solve NLP tasks using theoretical and non-theoretical approaches. This yielded 95 papers which we comprehensively surveyed in this paper. Finally, we conclude that while the applications of TDA in NLP have greatly improved since 2013, there is still room for improvement, specifically in reducing the barrier to entry for non-TDA experts to apply it to their NLP tasks.
Acknowledgments
Adaku would like to thank Dr. Charlie Dagli for his encouragement, reading the paper drafts, and providing invaluable recommendations.
Appendix A Case Studies of TDA for NLP Tasks
We showcase the utility of TDA in non-trivial NLP tasks by applying persistent homology and mapper using the Theoretical and Non-theoretical approaches. See below:
A.1 Theoretical Approaches.
(Semantic) Topological Space.
For the analysis, we use a subset of the C4 dataset that has watermarked and non-watermarked LLM-generated texts [^4] https://huggingface.co/datasets/acmc/watermarked_c4_dataset. Next, we use the $ripser$ python package to extract the topological features with persistent homology and $textdescriptives$ Python package [^5] https://github.com/HLasse/TextDescriptives?tab=readme-ov-file to extract the linguistic features.
Topological “Shape” of Words.
Using the same C4 dataset, we create Mapper diagrams using the $Mapper$ Python package [^6] https://github.com/scikit-tda/kepler-mapper/tree/master.
A.2 Non-Theoretical Approaches
Model Interpretation and Analysis.
We evaluate the Euclidean and Wasserstein distances between the embeddings of English-French translation text pairs [^7] https://huggingface.co/datasets/aircrypto/English-French-Translations (using the first 500 text pairs). For the Euclidean distance, we compare the embeddings from the English model [^8] https://huggingface.co/sentence-transformers/stsb-xlm-r-multilingual and the multilingual model [^9] https://huggingface.co/sentence-transformers/msmarco-roberta-base-v2. To calculate the Wasserstein distance, we first extract topological features from the embeddings, represent them as persistent diagrams, and then calculate the distance. Euclidean and Wasserstein distance range is [0, $\infty$).
Appendix B List of papers
See list of Theoretical and Non-Theoretical applications of TDA in NLP in Tables 4 and 5, respectively.
@@ Appendix B List of papers
Fitz et al. (2024) Topological “Shape” of Words Measures the topological complexity of Transformer-based hidden representations Persistent Homology
@@
Bouazzaoui et al. (2021) Topological “Shape” of Words Investigates the similarity between the topological shapes of the Tifinagh and Phoenician scripts Persistent Homology
@@
Sovdat (2016) distinguishing between languages TF-IDF Supervised Text Persistent Homology
@@
Short-PHD Wei et al. (2025) deepfake text detection (for short-text) Pre-trained (Transformers - CLS) Supervised Text Persistent Homology
Guilinger et al. (2025) deepfake text detection (for academic abstracts) Pre-trained (Transformers - CLS) Supervised Text Persistent Homology
@@
Alexander and Wang (2023) Hate speech, Misinformation & Psychiatric disorder classification Pre-trained (Transformers - Hidden) Supervised Text Mapper
Persistent BERT Compression and Explainability (PBCE) Balderas et al. (2025) Model compression Pre-trained (Transformers - Hidden) Unsupervised Text Persistent Homology
@@
Bazarova et al. (2025) LLM Hallucination detection Pre-trained (Transformers - Attention) Supervised Text Persistent Homology
@@
Sakurai et al. (2025) Authorship attribution of Japanese texts Pre-trained (Transformers - Attention) Supervised Text Persistent Homology
@@
Proskura and Zaytsev (2024) estimation of weights for ensembles of classification models Pre-trained (Transformers - Attention, CLS) Supervised Text Persistent Homology
Snopov and Golubinskiy (2024) vulnerability detection in code Pre-trained (Transformers - Attention) Supervised Text Persistent Homology
@@
Vu et al. (2025) multi-modal adversarial robustness assessment Multi-Modal Supervised Image & Text Persistent Homology